OpenStudy (anonymous):

A piece of tile is 4 1/2 inches long and 3 1/4 inches wide. What is the area of the tile? A. 7 3/4 square inches B. 12 square inches C. 12 1/8 square inches D. 14 5/8 square inches

OpenStudy (anonymous):

@jigglypuff314 @sleepyhead314 @hhelpplzzzz

OpenStudy (anonymous):

plz help

OpenStudy (raven7266):

D

OpenStudy (anonymous):

can you plz explain

OpenStudy (anonymous):

will fan and medal

OpenStudy (raven7266):

\[(9/2)x(13x4)=117/8 or 14(5/8)\]

OpenStudy (anonymous):

thanks

OpenStudy (raven7266):

I mean 13/4

OpenStudy (anonymous):

wait what?

OpenStudy (texaschic101):

4 1/2 * 3 1/4 = 9/2 * 13/4 = 117/8 = 14 5/8

OpenStudy (anonymous):

wait @raven7266 said 13/4

OpenStudy (texaschic101):

she made a typo...it was supposed to be 13/4 not 13 * 4

OpenStudy (anonymous):

oh ok :P i wish i could give you a medal to

OpenStudy (raven7266):

haha no I mistype the thing I wrote earlier what I'm trying to say is I wrote 13x4 instead of 13/4.. @texaschic 101 thanks for better explanation

OpenStudy (triciaal):

Area of a rectangle is length times width Here 4 1/2 * 3 1/4 Unit of measure for area is square units. Here inches

OpenStudy (texaschic101):

you gave the right person a medal

OpenStudy (anonymous):

i already have the answer @triciaal :3 i dont need help anymore